The difference of two numbers is 13, Five times the smaller number is one more than double the larger number. Find the two numbers.
Answer by mananth(16949)

let the numbers be x & y
x-y =13..........1
5y=2x+1
5y-2x=1............2
multiply (1) by 2
2x-2y=26
add to (2)
3y=27
y=9
..
x-y=13
x-9=13
x=22
...
the numbers are 22,9